Skip to end of metadata
Go to start of metadata

We just realized that 13b was still "Tee" wired into the 3.3m sonic, where it should have been changed during the array2->array3 shift to the sonic now at 4.8m.  I just did this at 1136am.  Unfortunately, Gordon will have to do his magic to fix this for the first 3 days of array3.  Sorry...